# Reinforcement Unlearning

### v2

@@ 3.2. Decremental RL-based Method
The implementation of this method involves two main steps. The first one is the exploration of the unlearning environment $\mathcal{M}_{u}$. Initially, the agent is allowed to explore the unlearning environment, collecting experience samples specific to that environment. The nature of this exploration depends on the scenario. For instance, in the grid-world setting, the agent might traverse the unlearning grid using a random walk for a predefined number of steps. In the aircraft-landing scenario, it could involve the airplane making random landings within the unlearning environment. The second step is fine-tuning the agent. Following the exploration phase, the agent is fine-tuned using the collected experience samples. This fine-tuning process employs a newly defined loss function (Eq. 5) to update the policy $\pi^{*}$ with the experience samples accumulated in the first step. This loss function is carefully designed to ensure that the agent’s performance within the unlearning environment $\mathcal{M}_{u}$ degrades while preserving its performance in other environments. Essentially, it guides the agent to unlearn the knowledge associated with the unlearning environment.
@@
Figure 1. The schematic diagram of the decremental reinforcement learning-based method
@@
To accomplish the aim of unlearning, we establish an optimization objective to guide the unlearning process (Eqs. 3 and 4). We also introduce a new loss function (Eq. 5) that will be used to update the agent. This loss function is designed to minimize the influence of the previously learned knowledge and encourage the agent to modify its existing policy. By incorporating this loss function into the training procedure, we can steer the agent’s learning process towards unlearning the knowledge from the given environment.
(5) $\mathcal{L}_{u}=\mathbb{E}_{s\sim\mathcal{S}_{u}}[||Q_{\pi^{\prime}}(s)||_{\infty}]+\mathbb{E}_{s\not\sim\mathcal{S}_{u}}[||Q_{\pi^{\prime}}(s)-Q_{\pi}(s)||_{\infty}].$
@@
In Eq. 5, the first term encourages the new policy $\pi^{\prime}$ to work deficiently in the unlearning environment $\mathcal{M}_{u}$, while the second term drives the new policy $\pi^{\prime}$ to have the same performance as the current policy $\pi$ in other environments. Notably, the two terms in Eq. 5 have favorable properties. The first term directs an agent to search and attempt different policies to sufficiently explore the state space of environment $\mathcal{M}_{u}$. Thus, $\mathcal{M}_{u}$ can be adequately unlearned. This property is particularly useful when $\mathcal{M}_{u}$ is a sparse reward setting, i.e., the reward is 00 in most of the states in $\mathcal{S}_{u}$. The second term motivates an agent to modify policies, instead of randomly changing its behavior. This property ensures that the agent performs consistently in those states which are not in $\mathcal{S}_{u}$. It is essential to highlight that the accurate computation of the second term is infeasible due to its involvement with all states except $\mathcal{S}_{u}$. Thus, during implementation, we uniformly select a consistent set of states across all environments, excluding $\mathcal{M}_{u}$. This approach is taken to mitigate computational burden and ensure a balanced impact on performance preservation across the remaining environments.
@@
The decremental RL-based method is formally described as follows. First, the agent explores the unlearning environment using a random policy. This policy ensures that the agent takes each action with equal probability. Adopting a random policy ensures a comprehensive exploration of the unlearning environment. If the agent were to use a well-established policy from its training, there is a risk of swift achievement of the target, potentially resulting in insufficient collection of experience samples within the unlearning environment. This strategic use of a random policy enables the effective collection of diverse experiences crucial for the unlearning process. For instance, in the grid world setting, where the agent has four possible actions (moving up, down, left, and right), the random policy dictates that, in each grid (i.e., state), the agent has an equal likelihood of selecting any of the four directions. As the agent progresses in its exploration, in each time step $t$, after taking an action $a_{t}$, it receives a corresponding reward $r_{t}$. Coupled with the current state $s_{t}$ and the subsequent state $s_{t+1}$, the agent effectively creates an experience sample in the form of $(s_{t},a_{t},r_{t},s_{t+1})$. In our method, the agent is directed to explore the unlearning environment for a specified number of steps, denoted as $m$. Consequently, the agent accumulates a total of $m$ experience samples: $(s_{1},a_{1},r_{1},s_{2}),\ldots,(s_{m},a_{m},r_{m},s_{m+1})$. In the second step, the agent employs these collected experience samples to fine-tune its current optimal policy $\pi^{*}$ to a new policy $\pi^{\prime}$ by minimizing the custom loss function defined in Eq. 5.
@@
For the first term in Eq. 5, let $y_{\pi}=r(s,\pi(s))+\gamma max_{\pi(s^{\prime})}Q(s^{\prime},\pi(s^{\prime}))$ and $\delta_{\pi}=y_{\pi}-Q(s,\pi(s))$, where $s^{\prime}$ denotes the next state. Then, the loss function in Eq. 2 can be rewritten as:
(6) $\mathcal{L}=\mathbb{E}_{s\sim\mathcal{S}}[\delta_{\pi}].$
@@
(7) $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}-\delta_{\pi^{\prime}}|]\leq\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}|]+\mathbb{E}_{s\sim\mathcal{S}_{u}}[|\delta_{\pi^{\prime}}|].$
As the convergence of the learning on the loss function in Eq. 6 has been both theoretically and empirically proven (Mnih et al., 2013; Tosatto et al., 2017; Fan et al., 2020), we can also conclude the convergence of $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|\delta_{\pi^{\prime}}|]$ in Eq. 7. For the first term in the right of Eq. 7, $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}|]$, as it is computed by accumulating the previously collected discounted rewards (Eq. 1 and 2), the term converges if the rewards are bounded. The reward bound can be acquired by proper definition, i.e., $r\in[-R_{max},R_{max}]$. Thus, as both $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}|]$ and $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|\delta_{\pi^{\prime}}|]$ converge, $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}-\delta_{\pi^{\prime}}|]$ also converges, i.e., $\mathbb{E}_{s\sim\mathcal{S}_{u}}[||Q_{\pi^{\prime}}(s)||_{\infty}]$ converges.
@@
Theorem 1 (Error Propagation (Tosatto et al., 2017)).
@@
Theorem 1 provides evidence that the disparity between the learned Q-function and the optimal Q-function diminishes as the learning process advances. This reduction signifies convergence, given that the Q-function, denoted as $Q$, remains uniformly bounded by $\frac{R_{max}}{1-\gamma}$ for any policy $\pi$ (Tosatto et al., 2017). Consequently, if the number of learning iterations is sufficiently large, the method converges.
@@
3.3. Environment Poisoning-based Method
@@
This method is implemented by modifying the unlearning environment itself. This modification can include various changes, i.e., the poisoning actions, such as altering the layout in the grid world scenario by adding or removing obstacles and repositioning targets within the environment. After these changes are introduced, the agent is updated in this modified environment. This method aims to influence the agent’s policy learning by creating a situation where its previously learned knowledge becomes less effective, particularly within the context of the unlearning environment.
@@
Specifically, this method consists of three distinct steps. Firstly, we apply a random poisoning strategy to alter the transition function of the unlearning environment. Secondly, the agent learns a new policy in this modified environment. Lastly, based on the agent’s learned policy, we update the poisoning strategy and re-poison the unlearning environment. These three steps are iteratively repeated until a predetermined number of poisoning epochs is reached, effectively refining the unlearning process. By employing this iterative approach, the method enhances the agent’s ability to unlearn specific experiences associated with the targeted unlearning environment. The schematic diagram of this method is presented in Figure 2, illustrating that the targeted unlearning environment $\mathcal{M}_{u}$ is altered to a new one $\mathcal{M}^{\prime}_{u}$ with strategically introduced perturbations, e.g., adding fake obstacles, and the agent is retrained in this poisoned environment to learn a new policy $\pi^{\prime}$.
Figure 2. The schematic diagram of the environment poisoning-based method
@@
Let us consider the learned policy as $\pi^{*}$, which is regarded as the optimal policy. To refine the agent’s policy, we manipulate the given environment $\mathcal{M}_{u}$. The manipulation of $\mathcal{M}_{u}$ involves poisoning the transition function $\mathcal{T}_{u}(s^{\prime}|s,a)$ (Xu et al., 2022), where $\mathcal{M}_{u}=\langle\mathcal{S}_{u},\mathcal{A}_{u},\mathcal{T}_{u},r\rangle$. We introduce a poisoned transition function denoted as $\hat{\mathcal{T}}_{u}(\hat{s}^{\prime}|s,a)$. After the agent takes action $a$ in state $s$, instead of observing the intended state $s^{\prime}$, it will observe the manipulated state $\hat{s}^{\prime}$. The challenge now lies in determining the appropriate state $\hat{s}^{\prime}$, which can mislead the agent’s learning process. To address this, we define a new learning environment for poisoning, denoted as $\mathcal{M}_{p}=\langle\mathcal{O},\mathcal{G},\mathcal{P},\mathcal{R}\rangle$. By constructing this new environment, we can manipulate the state transitions experienced by the agent, thus guiding its learning process in a desired manner.
@@
  • $\mathcal{O}$ denotes the set of poisoning states. Each state $\pi_{i}\in\mathcal{O}$ is the policy used by the agent during the $i$-th poisoning epoch.
@@
  • $\mathcal{G}$ is the set of poisoning actions. A poisoning action $g\in\mathcal{G}$ signifies a modification made to the transition function of the unlearning environment $\mathcal{T}_{u}$. This modification determines which state should be presented to the agent as the new state.
@@
  • $\mathcal{P}:\mathcal{O}\times\mathcal{G}\times\mathcal{O}\rightarrow[0,1]$ defines the poisoning state transition. It describes how the agent adjusts its policy in response to poisoning actions. Specifically, $\mathcal{P}(\pi^{\prime}|\pi,g)$ is the probability of the agent transitioning from policy $\pi$ to policy $\pi^{\prime}$ when the unlearning environment’s transition function is modified by $g$.
@@
    (8) $\mathcal{R}_{i}:=\lambda_{1}\Delta(\pi_{i}(s_{i})||\pi^{\prime}(s_{i}))+\lambda_{2}\sum_{s\not\sim\mathcal{S}_{u}}\sum_{a}\pi_{i}(s,a)r(s,a).$
    $\mathcal{R}_{i}$ represents the reward received during the $i$-th poisoning epoch. The term $\Delta(\pi_{i}(s_{i})||\pi^{\prime}(s_{i}))$ indicates the difference between $\pi_{i}(s_{i})$ and $\pi^{\prime}(s_{i})$. Here, $\pi_{i}(s_{i})$ represents the probability distribution over the available actions in state $s_{i}$ under policy $\pi_{i}$. This difference can be measured using either KL-divergence or cosine similarity. Coefficients $\lambda_{1}$ and $\lambda_{2}$ are introduced to balance the two terms. Note that precisely computing the second term is computationally infeasible due to the involvement of states from all environments except $\mathcal{M}_{u}$. Thus, similar to the decremental RL-based method, in the implementation of the poisoning-based method, a uniform selection of states across all the environments, except $\mathcal{M}_{u}$, is performed in each poisoning epoch.
@@
The proposed poisoning-based method is outlined in Algorithm 1. In each poisoning epoch $i$, we take the first step by choosing a poisoning action $g_{i}$ to modify the transition function of the unlearning environment (Lines 1-3). The selection process can be implemented using an $\epsilon$-greedy strategy, where the best action is chosen with a probability of $1-\epsilon$, and the remaining actions are chosen uniformly with a probability of $\frac{\epsilon}{|\mathcal{G}|-1}$. Here, the “best action” denotes the action that results in the highest Q-value, signifying the maximum expected future reward. This criterion guides the agent to identify the most effective way to modify the unlearning environment, ensuring optimal adjustments based on the expected outcomes.
Next, the agent performs the second step by learning a new policy $\pi_{i}$ in the altered environment (Line 4). This learning phase can be performed using any deep reinforcement learning algorithm, such as deep Q-learning (Mnih et al., 2013). Once $\pi_{i}$ is learned, we execute the third step by evaluating the reward using Eq. 8 and updating the poisoning strategy using samples $(\pi_{i-1},g_{i},\pi_{i},\mathcal{R}_{i})$ from batch $\mathcal{B}$ (Lines 5 and 6). The update process is carried out using the DDPG algorithm (Lillicrap et al., 2016). After all poisoning epochs are completed, we obtain a refined policy $\hat{\pi}$ (Line 8), which allows the agent to perform poorly in the unlearning environment $\mathcal{M}_{u}$, while maintaining satisfactory performance in other environments.
Convergence Analysis of the Method.  Algorithm 1 represents a deep RL algorithm employed by the model owner. This algorithm utilizes the agent’s policies as states, the modification of the unlearning environment as actions, and the performance of the agent’s policies in the environments as rewards. The interaction between the algorithm and the agent’s learning process is illustrated in Figure 3. In this figure, the model owner is engaged in learning how to poison the unlearning environment $\mathcal{M}_{u}$, while the agent is concurrently learning within the poisoned environment $\mathcal{M}^{\prime}_{u}$.
Figure 3. The interaction between agent learning and model owner poisoning. The owner observes the poisoning states about the unlearning environment $\mathcal{M}_{u}$, takes poisoning actions against $\mathcal{M}_{u}$, and receives poisoning rewards. The agent, on the other hand, is learning in the poisoned environment.
@@
where $d_{0}$ is the initial state distribution and $\mu_{\pi}>0$ for each state $s$. Here, $\mu_{\pi}$ satisfies the following Bellman flow constraints (Rakhsha et al., 2021):
(9) $\mu_{\pi}=(1-\gamma)d_{0}+\gamma\sum_{s^{\prime}}\mathcal{T}(s^{\prime}|\pi(s^{\prime}),s)\mu_{\pi(s^{\prime})}.$
@@
3.4. Comparison of the Two Methods
We compare the two methods in their ability of overcoming the “over-unlearning” and “catastrophic forgetting” issues.
In the context of a set of environments $\{\mathcal{M}_{1},...,\mathcal{M}_{n}\}$ and an unlearning environment $\mathcal{M}_{u}$, the objective of the decremental RL-based method is to minimize the agent’s return specifically in $\mathcal{M}_{u}$ while ensuring that the return in the remaining environments remains unaffected. In contrast, the poisoning-based method aims to develop a strategy to modify the unlearning environment $\mathcal{M}_{u}$ into $\mathcal{M}^{\prime}_{u}$ and subsequently fine-tune the agent within this modified environment. The objective is to ensure that the agent adapts well to $\mathcal{M}^{\prime}_{u}$ while exhibiting poor performance in the original unlearning environment $\mathcal{M}_{u}$. Thus, the key distinction between the two methods lies in their approaches. The decremental RL-based method focuses on erasing the knowledge acquired by the agent in $\mathcal{M}_{u}$, while the poisoning-based method encourages the agent to learn in a new but altered environment, i.e., $\mathcal{M}^{\prime}_{u}$.

### v5

@@ III-B Decremental RL-based Method
The implementation of this method involves two main steps. The first one is the exploration of the unlearning environment $\mathcal{M}_{u}$. Initially, the agent is allowed to explore the unlearning environment, collecting experience samples specific to that environment. The nature of this exploration depends on the scenario. For instance, in the grid-world setting, the agent might traverse the unlearning grid using a random walk for a predefined number of steps. The second step is fine-tuning the agent using the collected experience samples. This fine-tuning process employs a newly defined loss function (Eq. 5) to update the policy $\pi^{*}$ with the experience samples accumulated in the first step. This loss function is carefully designed to ensure that the agent’s performance within the unlearning environment $\mathcal{M}_{u}$ degrades while preserving its performance in other environments. Essentially, it guides the agent to unlearn the knowledge associated with the unlearning environment. The schematic diagram of this method is depicted in Figure 1, where the agent’s policy is updated from the optimal policy $\pi^{*}$ to a deteriorative policy $\pi^{\prime}$.
@@
Figure 1: The schematic diagram of the decremental reinforcement learning-based method
@@
To accomplish the aim of unlearning, we establish an optimization objective to guide the unlearning process (Eqs. 3 and 4). We also introduce a new loss function (Eq. 5) that will be used to update the agent. This loss function is designed to minimize the influence of the previously learned knowledge and encourage the agent to modify its existing policy. With this loss function, we can steer the agent’s learning process towards unlearning the knowledge from the given environment.
$\mathcal{L}_{u}=\mathbb{E}_{s\sim\mathcal{S}_{u}}[||Q_{\pi^{\prime}}(s)||_{\infty}]+\mathbb{E}_{s\not\sim\mathcal{S}_{u}}[||Q_{\pi^{\prime}}(s)-Q_{\pi}(s)||_{\infty}].$ (5)
@@
In Eq. 5, the first term encourages the new policy $\pi^{\prime}$ to work deficiently in the unlearning environment $\mathcal{M}_{u}$, while the second term drives the new policy $\pi^{\prime}$ to have the same performance as the current policy $\pi$ in other environments. Notably, the two terms in Eq. 5 have favorable properties. The first term directs an agent to search and attempt different policies to sufficiently explore the state space of environment $\mathcal{M}_{u}$. Thus, $\mathcal{M}_{u}$ can be adequately unlearned. This property is particularly useful when $\mathcal{M}_{u}$ is a sparse reward setting, i.e., the reward is 00 in most of the states in $\mathcal{S}_{u}$. The second term motivates an agent to strategically modify policies. This property ensures that the agent performs consistently in those states which are not in $\mathcal{S}_{u}$. Note that the accurate computation of the second term is infeasible due to its involvement with all states except $\mathcal{S}_{u}$. Thus, during implementation, we uniformly select a consistent set of states across all environments, excluding $\mathcal{M}_{u}$. This approach is taken to mitigate computational burden and ensure a balanced impact on performance preservation across the remaining environments.
@@
The decremental RL-based method is formally described as follows. First, the agent explores the unlearning environment using a random policy. Here, a random policy, defined by the agent owner, refers to a policy where the agent selects actions uniformly at random from the set of all possible actions available in any given state. This means that the probability of selecting any action $a$ from the set of possible actions $\mathcal{A}$ is $1/(|\mathcal{A}|)$, where $|\mathcal{A}|$ is the number of available actions in that state. For instance, in the grid world setting, where the agent has four possible actions (moving up, down, left, and right), the random policy dictates that, in each grid (i.e., state), the agent has an equal likelihood of selecting any of the four directions. This policy ensures that the agent does not leverage any previously learned strategies specific to the unlearning environment. If the agent were to use a well-established policy from its training, there is a risk of swift achievement of the target, resulting in insufficient collection of experience samples within the unlearning environment. During its exploration, in each time step $t$, after taking an action $a_{t}$, the agent receives a corresponding reward $r_{t}$. Coupled with the current state $s_{t}$ and the subsequent state $s_{t+1}$, the agent effectively creates an experience sample in the form of $(s_{t},a_{t},r_{t},s_{t+1})$. When the agent is directed to explore the unlearning environment for a specified number of steps, denoted as $m$, the agent accumulates a total of $m$ experience samples: $(s_{1},a_{1},r_{1},s_{2}),\ldots,(s_{m},a_{m},r_{m},s_{m+1})$. In the second step, the agent employs these collected experience samples to fine-tune its current optimal policy $\pi^{*}$ to a new policy $\pi^{\prime}$ by minimizing the custom loss function defined in Eq. 5.
@@
Let $y_{\pi}=r(s,\pi(s))+\gamma max_{\pi(s^{\prime})}Q(s^{\prime},\pi(s^{\prime}))$ and $\delta_{\pi}=y_{\pi}-Q(s,\pi(s))$. Then, Eq. 2 can be rewritten as:
$\mathcal{L}=\mathbb{E}_{s\sim\mathcal{S}}[\delta_{\pi}].$ (6)
@@
$\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}-\delta_{\pi^{\prime}}|]\leq\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}|]+\mathbb{E}_{s\sim\mathcal{S}_{u}}[|\delta_{\pi^{\prime}}|].$ (7)
As the convergence of the learning on the loss function in Eq. 6 has been both theoretically and empirically proven [27, 38, 8], we can also conclude the convergence of $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|\delta_{\pi^{\prime}}|]$ in Eq. 7. For the term $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}|]$ of Eq. 7, as it is computed by accumulating the previously collected discounted rewards (Eq. 1 and 2), the term converges if the rewards are bounded. The reward bound can be acquired by proper definition, i.e., $r\in[-R_{max},R_{max}]$. Thus, as both $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}|]$ and $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|\delta_{\pi^{\prime}}|]$ converge, $\mathbb{E}_{s\sim\mathcal{S}_{u}}[|y_{\pi^{\prime}}-\delta_{\pi^{\prime}}|]$ also converges, i.e., $\mathbb{E}_{s\sim\mathcal{S}_{u}}[||Q_{\pi^{\prime}}(s)||_{\infty}]$ converges.
@@
Theorem 1 (Error Propagation [38]).
@@
Theorem 1 provides evidence that the disparity between the learned Q-function and the optimal Q-function diminishes as the learning process advances. This reduction signifies convergence, given that the Q-function, denoted as $Q$, remains uniformly bounded by $\frac{R_{max}}{1-\gamma}$ for any policy $\pi$ [38]. Consequently, if the number of learning iterations is sufficiently large, the method converges.
@@
Note that the first and second terms in Eq. 5 may not converge simultaneously due to their optimization for different objectives. However, their convergence contributes to the overall convergence of Eq. 5. When both terms converge, it indicates that the agent’s policy is gradually adjusting to achieve the desired objectives outlined by each term. As the agent updates its policy over time, these adjustments lead to a convergence of the Q-values towards their respective targets. This convergence implies that the agent’s policy becomes increasingly optimized to fulfill both objectives. Our experimental results, depicted in Figure 2, reveal that both terms converge gradually over time. Although they exhibit different rates of convergence, they still contribute to the overall convergence of Eq. 5.
Figure 2: Convergence of Eq. 5 and its two terms
III-C Environment Poisoning-based Method
@@
This method is implemented by modifying the unlearning environment itself. This modification can include various changes, i.e., the poisoning actions, such as altering the layout in the grid world scenario by adding or removing obstacles and repositioning targets within the environment. After these changes are introduced, the agent is updated in this modified environment. This method aims to influence the agent’s policy learning by creating a situation where its previously learned knowledge becomes less effective.
The poisoning process operates as a Markov decision process, resembling an RL problem. Here, the agent’s policies serve as states, its modifications to the unlearning environment act as actions, and the disparity between the current policy’s performance and that of the updated policy within the unlearning environment functions as the reward. Consequently, the poisoning strategy is updated using Eq. 2 as the loss function.
@@
Specifically, this method consists of three steps. Firstly, we apply a random poisoning strategy to alter the transition function of the unlearning environment. Secondly, the agent learns a new policy in this modified environment. Lastly, based on the agent’s learned policy, we update the poisoning strategy and re-poison the unlearning environment. These three steps are iteratively repeated until a predetermined number of poisoning epochs is reached. The schematic diagram is presented in Figure 3, illustrating that the unlearning environment $\mathcal{M}_{u}$ is altered to $\mathcal{M}^{\prime}_{u}$ with strategically introduced perturbations, e.g., adding fake obstacles, and the agent is retrained in this poisoned environment to learn a new policy $\pi^{\prime}$.
Figure 3: The schematic diagram of the environment poisoning-based method
@@
Let the learned policy be $\pi^{*}$, which is regarded as the optimal policy. To refine the agent’s policy, we manipulate the given environment $\mathcal{M}_{u}$, involving poisoning the transition function $\mathcal{T}_{u}(s^{\prime}|s,a)$ [45], where $\mathcal{M}_{u}=\langle\mathcal{S}_{u},\mathcal{A}_{u},\mathcal{T}_{u},r\rangle$. We introduce a poisoned transition function denoted as $\hat{\mathcal{T}}_{u}(\hat{s}^{\prime}|s,a)$. After the agent takes action $a$ in state $s$, instead of observing the intended state $s^{\prime}$, it will observe the manipulated state $\hat{s}^{\prime}$. The challenge now lies in determining the appropriate state $\hat{s}^{\prime}$, which can mislead the agent’s learning process. To address this, we define a new learning environment for poisoning, denoted as $\mathcal{M}_{p}=\langle\mathcal{O},\mathcal{G},\mathcal{P},\mathcal{R}\rangle$.
@@
  • $\mathcal{O}$ is the set of poisoning states. Each state $\pi_{i}\in\mathcal{O}$ is the policy used by the agent during the $i$-th poisoning epoch.
@@
  • $\mathcal{G}$ is the set of poisoning actions. A poisoning action $g\in\mathcal{G}$ is a modification made to the transition function of the unlearning environment. This modification signifies which state should be presented to the agent as the new state.
@@
  • $\mathcal{P}:\mathcal{O}\times\mathcal{G}\times\mathcal{O}\rightarrow[0,1]$ defines the poisoning state transition. It describes how the agent adjusts its policy in response to poisoning actions. Here, $\mathcal{P}(\pi^{\prime}|\pi,g)$ is the probability of the agent transitioning from policy $\pi$ to $\pi^{\prime}$ when the unlearning environment’s transition function is modified by $g$.
@@
    $\mathcal{R}_{i}:=\lambda_{1}\Delta(\pi_{i}(s_{i})||\pi^{\prime}(s_{i}))+\lambda_{2}\sum_{s\not\sim\mathcal{S}_{u}}\sum_{a}\pi_{i}(s,a)r(s,a).$ (8)
    $\mathcal{R}_{i}$ represents the reward received during the $i$-th poisoning epoch. The term $\Delta(\pi_{i}(s_{i})||\pi^{\prime}(s_{i}))$ indicates the difference between $\pi_{i}(s_{i})$ and $\pi^{\prime}(s_{i})$, which can be measured using KL-divergence. Here, $\pi_{i}(s_{i})$ is the probability distribution over the available actions in state $s_{i}$ under policy $\pi_{i}$, and $\lambda_{1}$ and $\lambda_{2}$ are introduced to balance the two terms. Precisely computing the second term is computationally infeasible due to the involvement of states from all environments except $\mathcal{M}_{u}$. Thus, similar to the decremental RL-based method, in the implementation of the poisoning-based method, a uniform selection of states across all the environments, except $\mathcal{M}_{u}$, is performed in each poisoning epoch.
@@
The poisoning-based method is outlined in Algorithm 1. In each poisoning epoch $i$, we take the first step by choosing a poisoning action $g_{i}$ to modify the transition function of the unlearning environment (Lines 1-3). The selection process can be implemented using an $\epsilon$-greedy strategy, where the best action, which results in the highest Q-value, is chosen with a probability of $1-\epsilon$, and the remaining actions are chosen uniformly with a probability of $\frac{\epsilon}{|\mathcal{G}|-1}$. Here, $\epsilon$ is typically set within the range of $0.1$ to $0.2$ to achieve a balance between exploration and exploitation.
Next, the agent performs the second step by learning a new policy $\pi_{i}$ in the altered environment (Line 4). This learning phase can be performed using any RL algorithms, such as deep Q-learning [27]. Once $\pi_{i}$ is learned, we execute the third step by evaluating the reward using Eq. 8 and updating the poisoning strategy using samples $(\pi_{i-1},g_{i},\pi_{i},\mathcal{R}_{i})$ from batch $\mathcal{B}$ (Lines 5 and 6). The update process is carried out using the DDPG algorithm [22]. After all poisoning epochs are completed, we obtain a refined policy $\hat{\pi}$ (Line 8), which allows the agent to perform poorly in the unlearning environment $\mathcal{M}_{u}$, while maintaining satisfactory performance in other environments.
Convergence Analysis of the Method.  Algorithm 1 represents a deep RL algorithm employed by the model owner. The interaction between the algorithm and the agent’s learning process is illustrated in Figure 4. In this figure, the model owner is engaged in learning how to poison the unlearning environment $\mathcal{M}_{u}$, while the agent is concurrently learning within the poisoned environment $\mathcal{M}^{\prime}_{u}$.
Figure 4: The interaction between agent learning and model owner poisoning. The owner observes the poisoning states about the unlearning environment $\mathcal{M}_{u}$, takes poisoning actions against $\mathcal{M}_{u}$, and receives poisoning rewards. The agent, on the other hand, is learning in the poisoned environment.
@@
where $d_{0}$ is the initial state distribution and $\mu_{\pi}>0$ for each state $s$. Here, $\mu_{\pi}$ satisfies the following Bellman flow constraints [30]:
$\mu_{\pi}=(1-\gamma)d_{0}+\gamma\sum_{s^{\prime}}\mathcal{T}(s^{\prime}|\pi(s^{\prime}),s)\mu_{\pi(s^{\prime})}.$ (9)
@@
III-D Discussion of the Two Methods